OpenStudy (bekkah323):

find the slope intercept form of the equation of the line passing through the point (-1,3) with slope -2/5

OpenStudy (anonymous):

I love these ^.^

OpenStudy (anonymous):

y=mx+c m=slope=-2/5 therfore y=-2/5+c to find c sub in the point therefore 3=-2/5(-1)+c therfore c=13/5 y=-2/5x+13/5

OpenStudy (anonymous):

Aw, it's y = mx+b!

OpenStudy (bekkah323):

which point do you sub? or does it matter?

OpenStudy (anonymous):

sub in the given point (-1,3)

OpenStudy (bekkah323):

got it thanks

OpenStudy (anonymous):

@bekkah323 fan me